Sandra is 5 times as old as her son Ross. 3 years ago she was 7 times his age. How old is Ross?
Answer by 303795(602)

Now
Ross' age = r
Sandra's age = 5r
Three years ago
Ross' age = r - 3
Sandra's age = 5r - 3
Sandra was 7 times Ross' age so
5r - 3 = 7(r - 3)
5r - 3 = 7r - 21
Subtract 5r from each side
5r - 3 - 5r = 7r - 21 - 5r
-3 = 2r - 21
Add 21 to each side
-3 + 21 = 2r - 21 + 21
18 = 2r
r = 9
So Ross is currently 9 and Sandra is 45. Three years ago he was 6 and she was 42.